# Quantum automorphism groups of 2-graphs

### v2

@@ 2.16. Quantum automorphism group of finite graphs
Recall the definition of finite, directed graph $\mathcal{G}=(V=\{v_{1},\ldots,v_{n}\},E=\{e_{1},\ldots,e_{m}\})$ without multiple edge and the CQG $S_{n}^{+}$. We shall call a finite, directed graph a $1$-graph to be consistent with higher rank graph terminology from now on.
@@
The quantum automorphism group of a graph $\mathcal{G}$ without multiple edges (to be denoted by ${\rm Aut}^{+}(\mathcal{G})$) is defined to be the quotient $S^{+}_{n}/(AD-DA)$, where $A=((q_{ij}))_{i,j=1,\ldots,n}$, and $D$ is the vertex matrix of $\mathcal{G}$. The coproduct on the generators is given by $\Delta(q_{ij})=\sum_{k=1}^{n}q_{ik}\otimes q_{kj}$.
For the classical automophism group ${\rm Aut}(\mathcal{G})$, the commutative CQG $C({\rm Aut}(\mathcal{G}))$ is generated by $q_{ij}$’s where $q_{ij}$ is a function on $S_{n}$ taking value $1$ on the permutation which sends $i$-th vertex to $j$-th vertex and takes the value zero on other elements of the group. It is a quantum subgroup of ${\rm Aut}^{+}(\mathcal{G})$. The surjective CQG morphism $\pi:C({\rm Aut}^{+}(\mathcal{G}))\rightarrow C({\rm Aut}(\mathcal{G}))$ sends the generators to generators.
Theorem 2.18.
(Lemma 3.1.1 of [Fulton]) The underlying $C^{\ast}$-algebra of the quantum automorphism group ${\rm Aut}^{+}(\mathcal{G})$ of a finite graph $\mathcal{G}$ with $m$ edges and $n$ vertices (without multiple edge) is the universal $C^{\ast}$-algebra generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$ satisfying the following relations:
@@ 3. Main Results
To define the quantum automorphism group of a $2$-graph, we shall define the notion of the quantum automorphism group of a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ first. Then we shall show that the quantum automorphism groups of two such equivalent triples are isomorphic. Then as the isomorphism class of a $2$-graph depends on the equivalence class of its defining triple, one can define the quantum automorphism group of a $2$-graph to be the quantum automorphism group of its defining triple unambiguously. To that end let $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ be a triple where we remind the readers that $\mathcal{G}_{i}=(V,E_{i})$ are $1$-graphs (without any multiple edge, source or sink) with common set of finitely many vertices $V$ and edge sets $E_{i}$’s for $i=1,2$ respectively; Let the vertex matrices of $\mathcal{G}_{i}$ be $A_{i}$ for $i=1,2$ respectively such that $A_{1}A_{2}=A_{2}A_{1}$. Given a CQG $\mathbb{G}$ having an action $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ given by $\alpha(\delta_{v_{i}})=\sum_{j}\delta_{v_{j}}\otimes q_{ji}$ for $q_{ji}\in C(\mathbb{G})$ such that $((q_{ij}))$ commutes with $A_{1}$ and $A_{2}$, there are corepresentations $\beta_{i}$ on $C(E_{i})$ given by
$\beta_{i}(\delta_{e^{i}_{\alpha}})=\sum_{\beta}\delta_{e^{i}_{\beta}}\otimes q_{t(e^{i}_{\beta})t(e^{i}_{\alpha})}q_{s(e^{i}_{\beta})s(e^{i}_{\alpha})}.$
Then one can consider tensor product corepresentations $\beta_{1}\otimes\beta_{2}$, $\beta_{2}\otimes\beta_{1}$ on $C(E_{1}\times E_{2})$ and $C(E_{2}\times E_{1})$ respectively. Now note that $C(E_{1}\star E_{2})$ is a linear subspace of $C(E_{1}\times E_{2})$ with a basis for $C(E_{1}\star E_{2})$ given by $\{\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}}:s(e^{1}_{\alpha})=t(e^{2}_{\beta})\}$. Similarly $C(E_{2}\star E_{1})$ is a linear subspace of $C(E_{2}\times E_{1})$.
@@
Now suppose $s(e^{1}_{\alpha})=t(e^{2}_{\beta})$ so that $\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}}\in C(E_{1}\star E_{2})$. Then for $(e^{1}_{\mu},e^{2}_{\nu})$ such that $s(e^{1}_{\mu})\neq t(e^{2}_{\nu})$, we have $q_{t(e^{1}_{\mu})t(e^{1}_{\alpha})}q_{s(e^{1}_{\mu})s(e^{1}_{\alpha})}q_{t(e^{2}_{\nu})t(e^{2}_{\beta})}q_{s(e^{2}_{\nu})s(e^{2}_{\beta})}=0$ which proves the lemma. $\Box$
@@
A CQG $\mathbb{G}$ is said to act on a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$, if there is a faithful action $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ given by $\alpha(\delta_{v_{i}})=\sum_{j}\delta_{v_{j}}\otimes q_{ji}$ such that
(i) The matrix $((q_{ij}))$ commutes with the vertex matrices of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ so that we have the corresponding linear maps $\beta$ and $\beta^{{\rm op}}$ on $C(E_{1}\star E_{2})$ and $C(E_{2}\star E_{1})$ respectively.
(ii) The following diagram commutes:
{tikzcd}
Generators and relations: Let $\mathbb{G}$ be a CQG acting on a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ so that (assume there are $n$-vertices) $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ is a faithful action given by $\alpha(\delta_{v_{i}})=\sum\delta_{v_{j}}\otimes q_{ji}$. Then $C(\mathbb{G})$ is generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$. Now $q_{ij}$’s satisfy all the relations of ${\rm Aut}^{+}(\mathcal{G}_{i})$ for $i=1,2$ as described in the Theorem 2.18. Let us figure out the relations imposed by the commutative diagram in the Definition (3.3). To that end we start with the easy observation that if $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$, then $\theta_{\ast}(\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}})=\delta_{e^{2}_{\mu}}\otimes\delta_{e^{1}_{\nu}}$. Let $(e^{1}_{\alpha},e^{2}_{\beta})\in E_{1}\star E_{2}$ such that $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ for some $e^{2}_{\mu},e^{1}_{\nu}\in E_{2}\star E_{1}$. Then
@@
Now let $\theta(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})=(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$ and observe that as $\theta$ is a bijection, if $(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})$ vary over all $E_{1}\star E_{2}$, $(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$ vary over all of $E_{2}\star E_{1}$ so that
@@
(3.2) $=\sum_{(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})\in E_{2}\star E_{1}}\delta_{e^{2}_{\mu^{\prime}}}\otimes\delta_{e^{1}_{\nu^{\prime}}}\otimes q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}$
Comparing coefficients, we get for all $(e^{1}_{\alpha},e^{2}_{\beta})\in E_{1}\star E_{2}$ and all $(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})\in E_{2}\star E_{1}$ with $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ and $\theta(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})=(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$,
@@
Observation: Recall that $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ and $\theta(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})=(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$Rightarrow $t(e^{2}_{\mu^{\prime}})=t(e^{1}_{\alpha^{\prime}})$, $t(e^{2}_{\mu})=t(e^{1}_{\alpha})$, $s(e^{1}_{\nu^{\prime}})=s(e^{2}_{\beta^{\prime}})$ and $s(e^{1}_{\nu})=s(e^{2}_{\beta})$. Therefore the relations (3.3) have a few equivalent variations as $q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}=q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}$ and $q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}=q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}$.
@@
The quantum automorphism group of a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ (to be denoted by ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$) is defined to be the universal CQG in the category of CGQ’s acting on the triple in the sense of Definition 3.3.
@@
Let $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ be a triple such that $\mathcal{G}_{1},\mathcal{G}_{2}$ have a common set of vertices $V$ with $n$-points and both the $1$-graphs do not admit any multiple edge, source or sinks. Then ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ exists. The underlying $C^{\ast}$-algebra of ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ is the universal $C^{\ast}$-algebra generated by $n^{2}$-elements $q_{ij}$’s such that
(i) The elements $q_{ij}$’s satisfy the relations of ${\rm Aut}^{+}(\mathcal{G}_{i})$ for $i=1,2$.
(ii) The elements $q_{ij}$’s satisfy the relations (3.3).
The coproduct is given on the generators by $\Delta(q_{ij})=\sum_{k=1}^{n}q_{ik}\otimes q_{kj}$. If the usual basis for $C(V)$ is given by $\{\delta_{i}\}_{i=1}^{n}$, the coaction on $C(V)$ is given by $\alpha(\delta_{i})=\sum_{j=1}^{n}\delta_{j}\otimes q_{ji}$ for $i=1,\ldots,n$.
@@
$\Delta\Big{(}q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}q_{s(e^{2}_{\mu^{\prime}})s(e^{2}_{\mu})}q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}\Big{)}=\Delta\Big{(}q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}\Big{)}.$
@@
Now for $(j,i)\notin E_{2}$, $q_{t(e^{2}_{\mu^{\prime}})i}q_{s(e^{2}_{\mu^{\prime}})j}=0$. As none of the graphs $\mathcal{G}_{i}$ admits multiple edges, the last summation reduces to
@@
Since $\mathcal{G}_{1}$ is without source or sink, there is an edge $f^{1}_{\alpha^{\prime\prime}}\in E_{1}$ such that $s(f^{2}_{\mu^{\prime\prime}})=t(f^{1}_{\alpha^{\prime\prime}})$. In particular this implies that $(f^{2}_{\mu^{\prime\prime}},f^{1}_{\alpha^{\prime\prime}})\in E_{2}\star E_{1}$. Replacing $s(e^{2}_{\mu^{\prime}})$ by $t(e^{1}_{\nu^{\prime}})$ and using the fact that $q_{t(e^{1}_{\nu^{\prime}})t(f^{1}_{\alpha^{\prime\prime}})}q_{s(e^{1}_{\nu^{\prime}})k}=0$ for $k\neq s(f^{1}_{\alpha^{\prime\prime}})$, the last summation reduces to
@@
which can be shown to be equal to $\Delta\Big{(}q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}\Big{)}$. $\Box$

### v3

@@ 2.16. Quantum automorphism group of finite graphs
Recall the definition of a finite, directed graph $\mathcal{G}=(V=\{v_{1},\ldots,v_{n}\},E=\{e_{1},\ldots,e_{m}\})$ without multiple edges and the CQG $S_{n}^{+}$. We shall call a finite directed graph - a $1$-graph, to be consistent with higher rank graph terminology from now on.
@@
The quantum automorphism group of a graph $\mathcal{G}$ without multiple edges (to be denoted by ${\rm Aut}^{+}(\mathcal{G})$) is defined to be the quotient $S^{+}_{n}/\langle A\varepsilon-\varepsilon A\rangle$, where $A=((q_{ij}))_{i,j=1,\ldots,n}$, and $\varepsilon$ is the vertex matrix of $\mathcal{G}$. The coproduct on the generators is given by $\Delta(q_{ij})=\sum_{k=1}^{n}q_{ik}\otimes q_{kj}$.
For the classical automorphism group ${\rm Aut}(\mathcal{G})$, the commutative CQG $C({\rm Aut}(\mathcal{G}))$ is generated by $q_{ij}$’s where $q_{ij}$ is a function on $S_{n}$ taking value $1$ on the permutation which sends $i$-th vertex to $j$-th vertex and takes the value zero on other elements of the group. It is a quantum subgroup of ${\rm Aut}^{+}(\mathcal{G})$. The surjective CQG morphism $\pi:C({\rm Aut}^{+}(\mathcal{G}))\rightarrow C({\rm Aut}(\mathcal{G}))$ sends the generators to generators. We recall the following
Lemma 2.18 ([Fulton]*Lemma 3.1.1).
The underlying $C^{\ast}$-algebra of the quantum automorphism group ${\rm Aut}^{+}(\mathcal{G})$ of a finite graph $\mathcal{G}$ with $m$ edges and $n$ vertices (without multiple edges) is the universal $C^{\ast}$-algebra generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$ satisfying the following relations:
@@ 3. Main Results
To define the quantum automorphism group of a $2$-graph, we shall define the notion of the quantum automorphism group of a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ first. Then we shall show that the quantum automorphism groups of two such equivalent triples are isomorphic. Then as the isomorphism class of a $2$-graph depends on the equivalence class of its defining triple, one can define the quantum automorphism group of a $2$-graph to be the quantum automorphism group of its defining triple unambiguously. To that end, let $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ be a triple where we remind the readers that $\mathcal{G}_{l}=(V,E_{l})$ are $1$-graphs (bidirected and without any multiple edge, loops, or isolated vertices) with the common set of finitely many vertices $V$ and edge sets $E_{l}$’s for $l=1,2$ respectively; Let the vertex matrices of $\mathcal{G}_{l}$ be $\varepsilon_{l}$ for $l=1,2$ respectively such that $\varepsilon_{1}\varepsilon_{2}=\varepsilon_{2}\varepsilon_{1}$. Given a CQG $\mathbb{G}$ having an action $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ given by $\alpha(\delta_{v_{i}})=\sum_{j}\delta_{v_{j}}\otimes q_{ji}$ for $q_{ji}\in C(\mathbb{G})$ such that $((q_{ij}))$ commutes with $\varepsilon_{1}$ and $\varepsilon_{2}$, there are corepresentations $\beta_{l}$ on $C(E_{l})$ given by
$\beta_{l}(\delta_{e^{l}_{\alpha}})=\sum_{\beta}\delta_{e^{l}_{\beta}}\otimes q_{t(e^{l}_{\beta})t(e^{l}_{\alpha})}q_{s(e^{l}_{\beta})s(e^{l}_{\alpha})},$
where $e^{l}_{\alpha}$’s and $e^{l}_{\beta}$’s are edges. Then one can consider tensor product corepresentations $\beta_{1}\otimes\beta_{2}$ and $\beta_{2}\otimes\beta_{1}$ on $C(E_{1}\times E_{2})$ and $C(E_{2}\times E_{1})$ respectively. Now note that $C(E_{1}\star E_{2})$ is a linear subspace of $C(E_{1}\times E_{2})$ with a basis for $C(E_{1}\star E_{2})$ given by $\{\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}}:s(e^{1}_{\alpha})=t(e^{2}_{\beta})\}$. Similarly $C(E_{2}\star E_{1})$ is a linear subspace of $C(E_{2}\times E_{1})$.
@@
Now suppose $s(e^{1}_{\alpha})=t(e^{2}_{\beta})$ so that $\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}}\in C(E_{1}\star E_{2})$. Then for $(e^{1}_{\mu},e^{2}_{\nu})$ such that $s(e^{1}_{\mu})\neq t(e^{2}_{\nu})$, we have $q_{t(e^{1}_{\mu})t(e^{1}_{\alpha})}q_{s(e^{1}_{\mu})s(e^{1}_{\alpha})}q_{t(e^{2}_{\nu})t(e^{2}_{\beta})}q_{s(e^{2}_{\nu})s(e^{2}_{\beta})}=0$ which proves the lemma. ∎
@@
A CQG $\mathbb{G}$ is said to act on a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$, if there is an action $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ given by $\alpha(\delta_{v_{i}})=\sum_{j}\delta_{v_{j}}\otimes q_{ji}$ such that
  1. (i)
    The matrix $((q_{ij}))$ commutes with the vertex matrices of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ so that we have the corresponding linear maps $\beta$ and $\beta^{{\rm op}}$ on $C(E_{1}\star E_{2})$ and $C(E_{2}\star E_{1})$ respectively.
  2. (ii)
    The following diagram commutes:
    ${C(E_{1}\star E_{2})}$${C(E_{1}\star E_{2})\otimes C(\mathbb{G})}$${C(E_{2}\star E_{1})}$${C(E_{2}\star E_{1})\otimes C(\mathbb{G})}$$\scriptstyle{\beta}$$\scriptstyle{\theta_{\ast}}$$\scriptstyle{\theta_{\ast}\otimes{\rm id}}$$\scriptstyle{\beta^{\rm op}}$
The action is said to be faithful if $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ is faithful in the sense of Definition 2.12.
Generators and relations
Let $\mathbb{G}$ be a CQG acting on a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ so that (assume there are $n$-vertices) $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ is a faithful action given by $\alpha(\delta_{v_{i}})=\sum_{j}\delta_{v_{j}}\otimes q_{ji}$. Then $C(\mathbb{G})$ is generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$. Now $q_{ij}$’s satisfy all the relations of ${\rm Aut}^{+}(\mathcal{G}_{l})$ for $l=1,2$ as described in Theorem 2.18. Let us figure out the relations imposed by the commutative diagram in the Definition 3.3. We start with the easy observation that if $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$, then $\theta_{\ast}(\delta_{e^{1}_{\alpha}}\otimes\delta_{e^{2}_{\beta}})=\delta_{e^{2}_{\mu}}\otimes\delta_{e^{1}_{\nu}}$. Let $(e^{1}_{\alpha},e^{2}_{\beta})\in E_{1}\star E_{2}$ such that $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ for some $(e^{2}_{\mu},e^{1}_{\nu})\in E_{2}\star E_{1}$. Then
@@
Now let $\theta(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})=(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$ and observe that as $\theta$ is a bijection, if $(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})$ varies over all $E_{1}\star E_{2}$, $(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$ varies over all of $E_{2}\star E_{1}$ so that
@@
(3.2) $=\sum_{(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})\in E_{2}\star E_{1}}\delta_{e^{2}_{\mu^{\prime}}}\otimes\delta_{e^{1}_{\nu^{\prime}}}\otimes q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}.$
Comparing the coefficients, we get for all $(e^{1}_{\alpha},e^{2}_{\beta})\in E_{1}\star E_{2}$ and all $(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})\in E_{2}\star E_{1}$ with $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ and $\theta(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})=(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$,
@@
Observation
Recall that $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ and $\theta(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})=(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$ together imply $t(e^{2}_{\mu^{\prime}})=t(e^{1}_{\alpha^{\prime}})$, $t(e^{2}_{\mu})=t(e^{1}_{\alpha})$, $s(e^{1}_{\nu^{\prime}})=s(e^{2}_{\beta^{\prime}})$ and $s(e^{1}_{\nu})=s(e^{2}_{\beta})$. Therefore the relations (3.3) have a few equivalent variations as $q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}=q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}$ and $q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}=q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}$.
@@
The quantum automorphism group of a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ (to be denoted by ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$) is defined to be the universal CQG in the category of CGQ’s acting faithfully on the triple in the sense of Definition 3.3.
@@
Let $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ be a triple such that $\mathcal{G}_{1},\mathcal{G}_{2}$ have a common set of vertices $V$ with $n$-points and both the $1$-graphs are bidirected without any multiple edge, loops or isolated vertices. Then ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ exists. The underlying $C^{\ast}$-algebra of ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ is the universal $C^{\ast}$-algebra generated by $n^{2}$-elements $q_{ij}$’s such that
  1. (i)
    The elements $q_{ij}$’s satisfy the relations of ${\rm Aut}^{+}(\mathcal{G}_{l})$ for $l=1,2$.
  2. (ii)
    The elements $q_{ij}$’s satisfy the relations (3.3). The coproduct is given on the generators by $\Delta(q_{ij})=\sum_{k=1}^{n}q_{ik}\otimes q_{kj}$. If the usual basis for $C(V)$ is given by $\{\delta_{i}\}_{i=1}^{n}$, the coaction on $C(V)$ is given by $\alpha(\delta_{i})=\sum_{j=1}^{n}\delta_{j}\otimes q_{ji}$ for $i=1,\ldots,n$.
@@
$\Delta\left(q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}q_{s(e^{2}_{\mu^{\prime}})s(e^{2}_{\mu})}q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}\right)=\Delta\left(q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}\right).$
@@
Now for $(j,i)\notin E_{2}$, $q_{t(e^{2}_{\mu^{\prime}})i}q_{s(e^{2}_{\mu^{\prime}})j}=0$. As none of the graphs $\mathcal{G}_{l}$ admits multiple edges, the last summation reduces to
@@
Since $\mathcal{G}_{1}$ is bidirected and without isolated vertices, there is an edge $f^{1}_{\alpha^{\prime\prime}}\in E_{1}$ such that $s(f^{2}_{\mu^{\prime\prime}})=t(f^{1}_{\alpha^{\prime\prime}})$. In particular this implies that $(f^{2}_{\mu^{\prime\prime}},f^{1}_{\alpha^{\prime\prime}})\in E_{2}\star E_{1}$. Replacing $s(e^{2}_{\mu^{\prime}})$ by $t(e^{1}_{\nu^{\prime}})$ and using the fact that $q_{t(e^{1}_{\nu^{\prime}})t(f^{1}_{\alpha^{\prime\prime}})}q_{s(e^{1}_{\nu^{\prime}})k}=0$ for $k\neq s(f^{1}_{\alpha^{\prime\prime}})$, the last summation reduces to
@@
which can be shown to be equal to $\Delta\left(q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}\right)$. ∎
The next theorem will allow us to define the notion of the quantum automorphism group of a $2$-graph. The proof is more or less standard and hence omitted.